UT 20131013UT FLAO+PISCES DX

Observer: DMiller (LBTO),Arcidiacono,C, Boutsia, K, (INAF)
Telescope Operator:Becchetti G
SW Support:CBiddick
AO Support: DMiller
Telescope Support: KNewton
Instrument Support: Edwards,M

Summary

Link to INAF Observing log

Details

00:00 Carmelo powered on the WUnit, tested the Z-Stage, camera lens and communication to the AdSec. The WUnit power on from the AOSGUI failed (see below) but the power on from the WFS Control GUI worked fine.

Today at 0.00 I started the WFS system, and I noticed this error at the startup
at power on for AOS- Right panel:
THIS THE LOG
Sat Oct 12 23:54:58.226 2013 aos.info - WFS: starting power on sequence
Sat Oct 12 23:55:40.170 2013 aos.info - WFS: setting up BCUs...
Sat Oct 12 23:56:05.915 2013 aos.info - WFS: setting up flowerpot...
Sat Oct 12 23:57:06.631 2013 aos.WfsOn.failed - WfsOn[264] failed - Timeout waiting for picctrl.R.STATUS (-5001) TIMEOUT_ERROR

Just after I tried to turn on the system from the WFS control GUI and it seeems ok.
THIS THE LOG
Sat 12 Oct - 11:59:00 PM I'm repeating the operation from WFS Control GUI
Sat Oct 12 23:59:09.775 2013 aos.info - WFS: starting power on sequence
Sat Oct 12 23:59:09.792 2013 aos.info - WFS: setting up BCUs...
Sat Oct 12 23:59:09.806 2013 aos.info - WFS: setting up flowerpot...
Sat Oct 12 23:59:09.819 2013 aos.info - WFS: homing all movements
Sat Oct 12 23:59:58.344 2013 aos.info - Adsec 672a-right going SAFE and clean up.
Sat Oct 12 23:59:59.269 2013 aos.info - Updated variable adamhousekeeper.R.COIL_STATUS: 0
Sun Oct 13 00:00:01.012 2013 aos.info - No misfunctioning capacitive sensors found.
Sun Oct 13 00:00:01.025 2013 aos.info - No misfunctioning capacitive sensors found.
Sun Oct 13 00:00:01.637 2013 aos.info - Adsec 672a-right SAFE position reached.
Sun Oct 13 00:00:03.461 2013 aos.info - Updated variable adsecarb.R.FSM_STATE: Ready
Sun Oct 13 00:00:03.474 2013 aos.info - Updated variable AOARB.R.SL_READY: 0
Sun Oct 13 00:00:13.588 2013 aos.info - WFS: setup done.
Sun Oct 13 00:00:13.602 2013 aos.info - Updated variable wfsarb.R.FSM_STATE@M_FLAOWFS: Operating

01:20 Collect sky flats

01:35 I changed the Pisces Keyword to INSTRUME to match the change Craig had make to the Pisces Server. Before on every pisces_get_status they were receiving an Warning about not finding INSTRMT

01:40 Geno set pointing

01:44 Pisces Temperature 74.66

01:45 Preset and collimate.

02:50 5, 265 6, -6598 9 -468 10, 281 11 204

03:31 Change CCD47 Black Levels to 480, 470

-- DougMiller - 13 Oct 2013
Topic revision: r4 - 11 Nov 2013, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback